# FREA: Feasibility-Guided Generation of Safety-Critical Scenarios with Reasonable Adversariality

### v1

@@ B.2 Safe RL-based Setting
In this framework, each CBV acts as an RL agent tasked with attacking the AV while maintaining the AV’s feasibility. Previous works [5, 12] primarily focused on minimizing the distance between CBV and AV, often leading to unavoidable collisions, as discussed in Section 2. To mitigate this, we replace the collision reward with the goal-based adversarial reward, which encourages the CBV to reach a potential collision point with the AV. The configuration details are as follows:
@@
Reward. As previously discussed, we have replaced the collision reward with a goal-based reward. This adjustment encourages the CBV to navigate toward a potential collision point while avoiding collisions with other BVs. In practice, we set the local target for the AV as the potential collision point, which forms the overall reward function:
@@
PPO. This method employs an adversarial CBV policy based on PPO that aims to reach a potential collision point with the AV, utilizing the adversarial reward outlined in Eq. 19.
@@ C.1 Training Details about LFR
Figure 9: LFR visualization with 0.5m $d_{th}$ under various traffic scenarios. Red: AV. Blue: BV
@@ C.2 Application Details of LFR
During the feasibility-guided adversarial policy training for CBV, we encountered a significant challenge: the optimal feasible value function produces a single scalar value at each timestep. In scenarios with multiple CBVs, it is difficult to identify which CBV renders the AV’s operation unfeasible. To address this issue, we introduced a “pseudo” state for each CBV at every timestep. This “pseudo” state captures information relevant to both the AV and CBV from the perspective of the AV. Crucially, this “pseudo” state is utilized solely to assess the threat posed by each CBV to the AV at each timestep. By employing this method, we eliminate the need for the viewpoint transformation function $g(\cdot)$, previously described in Eqs. 6 and 8. Instead, the “pseudo” state is directly derived from the information available at each timestep, simplifying the process and enhancing the accuracy of LFR.
Appendix D Detail Explanation about Feasibility Metric
In Section 4.3, we introduce four key metrics to evaluate the feasibility of AV along collision trajectories induced by various CBV methods. Among these, the Infeasible Ratio (IR) and Infeasible Distance (ID) are novel metrics proposed in this paper. Their definitions and implications are detailed as follow:

### v2

@@ B.2 Safe RL-based Setting
In this framework, each CBV acts as an RL agent tasked with attacking the AV while maintaining the AV’s feasibility. Previous works [5, 12] primarily focused on minimizing the distance between CBV and AV, often leading to unavoidable collisions and unrealistic scenarios, as discussed in Section 2. To mitigate this, we replace the collision reward with the goal-based adversarial reward, which encourages the CBV to reach a potential conflict point with the AV. The configuration details are as follows:
@@
Reward. As previously discussed, we have replaced the collision reward with a goal-based reward. This adjustment encourages the CBV to navigate toward a potential conflict point while avoiding collisions with other BVs. In practice, we set the reference point in the global route of the AV as the potential conflict point, which forms the overall reward function:
@@
PPO. This method employs an adversarial CBV policy based on PPO that aims to reach a potential conflict point with the AV, utilizing the adversarial reward outlined in Eq. 18.
@@
KING. This method reimplements the optimization-based safety-critical scenario generation approach, KING, as presented in [5]. Given the differences between our general experimental setup and that of KING, we made several necessary adjustments. For example, KING’s experimental setup includes background vehicle counts limited to 1, 2, and 4, while we employ a continuous traffic flow with more than 10 vehicles. Additionally, the target speed for the AV in KING is set at $4\ m/s$, whereas we use a typical speed of $6\ m/s$. KING’s experiments are conducted across Town03 to Town06, while our work focuses on Town05 and Town02. Therefore, KING is not directly comparable with the mentioned baselines in experiments that quantify near-miss events and assess the generality of AV testing. As a result, we limit our comparison of KING with the baselines to AV’s feasibility experiments, which are less influenced by variations in traffic flow settings. Specifically, we adjust the target speed of the AV in KING from $4\ m/s$ to $6\ m/s$ and restrict our comparative analysis to results from Town05 to ensure fairness.
@@ C.1 Training Details about LFR
Figure 9: LFR visualization with 0.5m $d_{th}$ under various traffic scenarios.
@@ C.2 Application Details of LFR
During the feasibility-guided adversarial policy training for CBV, we encountered a significant challenge: the optimal feasible value function produces a single scalar value at each timestep. In scenarios with multiple CBVs, it is difficult to identify which CBV renders the AV’s operation unfeasible. To address this issue, we introduced a “pseudo” state for each CBV at every timestep. This “pseudo” state captures information relevant to both the AV and CBV from the perspective of the AV. Crucially, this “pseudo” state is utilized solely to assess the threat posed by each CBV to the AV at each timestep. By employing this method, we eliminate the need for the viewpoint transformation function $g(\cdot)$, previously described in Eqs. 6 and 9. Instead, the “pseudo” state is directly derived from the information available at each timestep, simplifying the process and enhancing the accuracy of LFR.
Appendix D Robustness of AV Pre-trained with Various CBV Methods
To explore the impact of reasonable adversarial scenarios on AV policy learning, we employed PPO [33] as the AV policy, which utilized the same forms of input and output as detailed for the CBV policy in Section B.2. By leveraging the AV’s reward function described in [25], we trained the PPO policy across various surrogate CBV methods. Subsequently, we conducted robustness evaluations on these well-trained AV policies, presenting the results in Table 5.
Table 5: Comparative performance of AVs pretrained with various CBV methods across different maps. Results are the average of 10 runs in “Scenario9” with varied seeds.
Town05 intersections Town02 intersections
Surr. CBV CBV CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑) CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑)
Standard Standard 11% 4m 19m 4% 68s 85 39% 8m 20m 18% 57s 70
PPO 17% 11m 17m 6% 66s 82 40% 11m 19m 15% 63s 70
FREA(Ours) 3% 6m 18m 1% 75s 89 40% 3m 18m 19% 56s 71
Standard FREA 39% 6m 17m 7% 66s 73 79% 3m 17m 35% 45s 51
PPO 36% 15m 21m 6% 66s 74 79% 4m 14m 37% 44s 51
FREA(Ours) 31% 12m 17m 5% 71s 76 73% 3m 20m 28% 51s 55
As shown in Table 5, our proposed FREA method, when used as the CBV policy for AV training, improves the Overall Score (OS) and reduces the Collision Rate (CR), thereby enhancing AV policy robustness. This improvement primarily stems from the reasonable adversarial CBV policy. In normal traffic, the rarity of safety-critical scenarios limits AVs’ ability to learn effective avoidance policies, especially for learning-based methods. Conversely, excessively adversarial policies like PPO often result in unavoidable collisions, hindering the development of correct avoidance policies since no policy can ensure AV safety. By increasing the frequency of AV-feasible near-miss events, our FREA method provides effective data for AV training, thus improving policy robustness.
Appendix E Detail Explanation about the Metrics
E.1 Near-miss Event Metrics
In traffic safety and vehicle dynamics, time-to-collision (TTC) and post-encroachment time (PET) are two commonly used metrics to quantify near-miss events[4]. Here are the formal definitions for these terms:
$\text{TTC}=\frac{d}{v_{r}}$, where $d$ represents the distance between the two vehicles or between a vehicle and an obstacle and $v_{r}$ is the relative velocity, which is the speed of one vehicle relative to the other or the obstacle.
$\text{PET}=t_{2}-t_{1}$, where $t_{1}$ is when the first vehicle passes through the intersection point and $t_{2}$ is when the second vehicle passes through the same point.
In our work, the TTC metric was calculated using a publicly available code library [30], and the PET metric was computed using the code provided in [4].
E.2 Feasibility Metrics
In Section 4.3, we introduce four key metrics to evaluate the feasibility of AV along collision trajectories induced by various CBV methods. Among these, the Infeasible Ratio (IR) and Infeasible Distance (ID) are novel metrics proposed in this paper. Their definitions and implications are detailed as follows: